# SurveyG: A Multi-Agent LLM Framework with Hierarchical Citation Graph for Automated Survey Generation

### v2

@@ Abstract.
Large language models (LLMs) are increasingly adopted for automating survey paper generation (Wang et al., [n. d.]; Liang et al., 2025; Yan et al., 2025; Su et al., 2025; Wen et al., 2025). Existing approaches typically extract content from a large collection of related papers and prompt LLMs to summarize them directly. However, such methods often overlook the structural relationships among papers, resulting in generated surveys that lack a coherent taxonomy and a deeper contextual understanding of research progress. To address these shortcomings, we propose SurveyG, an LLM-based agent framework that integrates hierarchical citation graph, where nodes denote research papers and edges capture both citation dependencies and semantic relatedness between their contents, thereby embedding structural and contextual knowledge into the survey generation process. The graph is organized into three layers: Foundation, Development, and Frontier, to capture the evolution of research from seminal works to incremental advances and emerging directions. By combining horizontal search within layers and vertical depth traversal across layers, the agent produces multi-level summaries, which are consolidated into a structured survey outline. A multi-agent validation stage then ensures consistency, coverage, and factual accuracy in generating the final survey. Experiments, including evaluations by human experts and LLM-as-a-judge, demonstrate that SurveyG outperforms state-of-the-art frameworks, producing surveys that are more comprehensive and better structured to the underlying knowledge taxonomy of a field.
Automated Survey Generation, Large Language Model, Literature Synthesis, Multi-agent, Hierarchical Graph Representation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Information extraction
@@ 1. Introduction
The exponential growth of research publications, particularly in rapidly evolving fields such as Artificial Intelligence (Huynh-The et al., 2023), has made it increasingly difficult for researchers to keep pace with new developments (Wang et al., [n. d.]; Wen et al., 2025). While survey papers serve as invaluable resources by synthesizing existing knowledge and identifying emerging trends, their manual construction is costly, time-consuming, and often unable to keep up with the overwhelming influx of literature (Liang et al., 2025). Although large language models (LLMs) offer promising text generation capabilities, they face critical limitations in handling massive reference sets, maintaining academic rigor, and providing up-to-date knowledge (Wu et al., 2025; Han et al., 2024). These challenges underscore the urgent need for an automated survey generation framework that can efficiently retrieve, organize, and synthesize literature into coherent, high-quality surveys tailored to users’ research interests.
Figure 1. Overview of the standard automated survey pipeline, which involves three core stages: (1) preparing relevant papers, (2) generating a structured outline that defines sections and subsections, and (3) composing the full survey.
@@
Some recent studies (Su et al., 2025; Wang et al., [n. d.]; Liang et al., 2025; Yan et al., 2025) have proposed autonomous survey generation frameworks based on user queries, following the basic pipeline illustrated in Figure 1. While these approaches represent promising progress, they exhibit two key limitations. Firstly, they neglect the relationships between papers, such as citation links, methodological connections, or subtopic dependencies, which are essential for understanding how works build upon one another, improve over foundational methods, and collectively shape research trends. Secondly, these frameworks employ a naive strategy for constructing structured outlines or full survey papers, simply concatenating summaries of individual papers. This not only exacerbates the long-context problem in LLMs but also fails to exploit the hierarchical organization of related works within subtopics.
@@
To address these limitations, we propose SurveyG. This autonomous survey generation system emphasizes knowledge representation of retrieved papers and employs hierarchical summarization to construct well-structured outlines, which are an essential component of high-quality surveys. In detail, we design an LLM-based multi-agent framework that represents knowledge using a hierarchical citation graph, where nodes correspond to papers and edges capture both citation relationships and semantic similarity. The graph is organized into three layers: Foundation, Development, and Frontier to reflect the progression of research from seminal contributions to incremental improvements and emerging directions. By combining horizontal searches within layers and vertical traversals across layers, our framework generates multi-aspect summaries that are subsequently consolidated into a structured survey outline via a multi-agent framework.
Figure 2. Evaluation of generated surveys across multiple metrics using LLM-as-a-judge, validated by human experts.
We evaluate SurveyG on 10 computer science topics from the SurGE benchmark (Su et al., 2025), comparing its survey generation performance with existing state-of-the-art frameworks (Liang et al., 2025; Yan et al., 2025; Wang et al., [n. d.]). Following prior work, we assess outlines along five dimensions: Coverage, Structure, Relevance, Synthesis, Critical Analysis. Details of the experimental setup are provided in Section 4. As shown in Figure 2, while baseline methods achieve reasonable performance in Relevance and partially in Structure and Coverage, they perform significantly worse in Synthesis and Critical Analysis, primarily due to their inability to model inter-paper relationships and the limitations imposed by long-context inputs.
In conclusion, this work presents three key contributions to automated survey generation. First, we introduce a hierarchical citation graph representation that models both citation and semantic relationships among papers. Second, we develop a graph-based traversal mechanism that operates across this hierarchy to produce diverse and multi-aspect summarizations, effectively capturing the methodological foundations, developmental trends, and frontier directions of a research field. Third, we design a multi-agent framework that combines retrieval-augmented generation (RAG) with pre-built hierarchical summaries as memory, allowing the system to automatically construct coherent and comprehensive survey drafts grounded in verifiable evidence.
2. Related works
2.1. Long-form Text Generation
LLMs have achieved remarkable progress, yet generating long-form, coherent, and logically structured documents remains a persistent challenge (Bai et al., 2023; Dong et al., 2023; Han et al., 2024). Recent works have explored different strategies to address the long-context problem. For example, Chain-of-Agents (Zhang et al., 2024) introduces a multi-agent collaboration framework where worker agents process segmented portions of text and a manager agent synthesizes them into coherent outputs, alleviating focus issues in long contexts. LongAlign (Bai et al., [n. d.]) proposes a recipe for long context alignment, combining instruction data construction, efficient batching, yielding strong gains on queries up to 100k tokens. Complementary to these, Xu et al. (Xu et al., 2023) systematically examine the trade-offs between retrieval-augmentation and context-window extension, showing that hybrid approaches can outperform both strategies alone. However, existing approaches often rely on raw reference texts, leading to inefficient retrieval, limited context utilization, and poor structural coherence in survey-like outputs.
2.2. Automatic Survey Generation
The automatic generation of literature reviews has been studied for over a decade, starting with multi-document summarization techniques that produced unstructured related work sections. Early systems, such as IBM Science Summarizer (Erera et al., 2019), focused on summarizing scientific articles, while more recent LLM-based methods like ChatCite (Li et al., 2024) and Susnjak et al. (Susnjak et al., 2025)’s domain-specific fine-tuning advanced the generation of comparative and knowledge-enriched reviews. Despite these advances, such methods primarily tackle summarization rather than the creation of fully structured survey papers. More recent systems, including AutoSurvey (Wang et al., [n. d.]), InteractiveSurvey (Wen et al., 2025), SurveyForge (Yan et al., 2025), and SurveyX (Liang et al., 2025), propose end-to-end pipelines integrating RAG, clustering, or multi-agent strategies to automate survey construction. These methods improve structural coherence and formatting consistency while scaling to long-form survey content. Nevertheless, most frameworks still restrict users to fixed input-output modes, overlooking relationships among papers and limiting interactivity, which often results in surveys that lack flexibility, relational awareness, and depth.
@@ 3. Methodology
We introduce SurveyG, an automated survey generation framework that operates in two main phases. The Preparation Phase involves retrieving relevant papers, summarizing their content, constructing a hierarchical citation graph, and extracting relationships by traversing the graph. The Generation Phase focuses on producing a structured outline and composing a complete survey by integrating instruction prompting within a multi-agent framework. An overview of SurveyG is illustrated in Figure 3.
@@ 3.1. Preparation Phase
We represent the relationships among papers using a hierarchical citation graph, where nodes correspond to academic papers and edges capture both citation links and semantic similarity, each weighted by a value $w$. Each node is further assigned to one of three layers: Foundation, Development, or Frontier, which reflect the role of the paper in the progression of research. Formally, the hierarchical citation graph is defined as $G=(V,E,L),$ where $V$ denotes the set of nodes (papers), $E\subseteq V\times V$ is the set of directed or undirected edges encoding citation or semantic relationships, and $L:V\to\{\text{Foundation},\text{Development},\text{Frontier}\}$ is a layer assignment function. For each node $v_{i}\in V$, we associate a corresponding document $d_{i}\in D$, where $D$ is a database storing the complete content of all papers. In addition, each node $v_{i}\in V$ is equipped with attributes that include a summarization of $d_{i}$ as well as metadata such as the paper’s title and publication year.
@@ 3.1.1. Searching Relevant Paper
where $\text{sim}(\cdot)$ denotes the cosine similarity between the text embedding vectors of the two papers. For computational efficiency, embeddings are derived solely from the abstract of each paper, which captures the core conceptual content while minimizing processing overhead.
To better leverage the key content of each paper for survey generation, every node $v_{i}$ is enriched with a summarization derived from its corresponding document $d_{i}$. We design specialized prompt templates tailored to different paper types (Liang et al., 2025)(e.g., surveys, methodological contributions, benchmarks, theoretical works). After this phase, we obtain a flat graph $\hat{G}=(V,E)$ that encompasses the papers relevant to the user’s topic along with their relationships. Each node $v_{i}\in V$ is associated with a set of attributes defined as
@@ 3.1.2. Knowledge Representation
To reflect the developmental progression of research within a topic, we assign each node in the flat graph $\hat{G}=(V,E)$ to one of three hierarchical layers via a layer assignment function
$L:V\to\{\text{Foundation},\text{Development},\text{Frontier}\}.$
(1) Foundation Layer. The foundation layer consists of seminal and high-impact works that form the intellectual backbone of the field. For each paper $p$, we define a trending score as
@@
You are a research analyst synthesizing papers on the topic [QUERY]. <think> Explain your reasoning for clustering papers into 2–3 subgroups based on methodology, contribution, or thematic focus. </think> For each subgroup, summarize the shared methodological approaches, thematic contributions, and provide a concise critique comparing the works. Finally, synthesize an overall perspective highlighting how these subgroups collectively operate in the field.
@@ 3.1.3. Traversal on Graph Strategy
In the vertical stage, we aim to model cross-layer dependencies. For each foundation paper, we perform a weighted breadth-first search (WBFS) over its citation paths, where traversal prioritizes semantically relevant nodes according to edge weights. The algorithmic details are provided in Algorithm 1, the full WBFS procedure and prompt design are described in Appendix A. Each resulting path aggregates the node attributes $A(v)$ for all $v$ encountered during the WBFS traversal and places them in the pathpath variable. We then apply hierarchical summarization across layers, exploiting temporal progression to mitigate long-context issues and extract key insights more effectively (Zhang et al., 2024). This process incrementally integrates knowledge from the Development and Frontier layers into path-specific summaries. After the summarization phase, SurveyG ultimately produces $K+N$ outputs with $N$ horizontal layer summaries and $K$ vertical path summaries.
In contrast to earlier frameworks (Wang et al., [n. d.]; Wen et al., 2025; Liang et al., 2025; Yan et al., 2025) that represent papers as isolated records in a flat database and depend exclusively on RAG-based retrieval, SurveyG organizes the literature within a hierarchical citation graph $G$. This representation integrates both citation and semantic connections among papers, allowing the system to capture the logical progression of research topics over time. By traversing this hierarchy, SurveyG produces a series of summarizations across multiple layers, effectively revealing methodological developments, evolutionary patterns, and current research frontiers. Such a design provides a more coherent and interpretable knowledge foundation for automated survey generation.
3.2. Generation Phase
We employ a multi-agent conversational framework (Wu et al., 2024) to guide the generation of survey papers. The system is composed of two complementary agents: a Writing Agent, equipped with memory (Sumers et al., 2023) initialized with $K+N$ summarizations from the graph traversal phase, and an Evaluation Agent, which leverages the internal reasoning capabilities of LLMs to provide diversity-oriented feedback. Through iterative interaction, the Writing Agent proposes structured content grounded in summarizations, while the Evaluation Agent critiques and refines these outputs to ensure coherence and balance. This cooperative setup enables the agents to jointly construct and improve survey papers by integrating both external evidence and internal reasoning. An overview of the overall generation process is provided in Algorithm 2.
@@ 3.2.1. Structured Outline Construction
The Writing Agent constructs an initial structured outline by grounding each section and subsection in the $K+N$ multi-aspect summarizations, ensuring both factual grounding and thematic coherence. The Evaluation Agent then reviews the draft, assessing logical flow and suggesting refinements without altering the overall structure. After one or two feedback iterations, the outline converges into a coherent and evidence-supported framework. Detailed prompting for both agents is provided in Figure 5 and Appendix C. The key innovation of SurveyG lies in its ability to manage long-context survey synthesis without concatenating all reference texts (Wang et al., [n. d.]; Liang et al., 2025) or relying on pre-existing human-written surveys (Yan et al., 2025). Instead, it leverages hierarchical summarization from the citation graph $G$ as structured knowledge injected into the Writing Agent.
Figure 5. Structured outline creation short version prompt.

### v3

# A Multi-Agent LLM Framework with Hierarchical Citation Graph for Automated Survey Generation

@@ Abstract.
Large language models (LLMs) are increasingly being adopted to automate survey paper generation. Existing approaches typically extract content from a large collection of related papers and prompt LLMs to summarize them directly. However, such methods often overlook the structural relationships among papers, resulting in generated surveys that lack a coherent taxonomy and a deeper contextual understanding of research progress. To address these shortcomings, we propose SurveyG, an LLM-based agent framework that integrates hierarchical citation graph, where nodes denote research papers and edges capture both citation dependencies and semantic relatedness between their contents, thereby embedding structural and contextual knowledge into the survey generation process. The graph is organized into three layers: Foundation, Development, and Frontier, to capture the evolution of research from seminal works to incremental advances and emerging directions. By combining horizontal search within layers and vertical depth traversal across layers, the agent produces multi-level summaries, which are consolidated into a structured survey outline. A multi-agent validation stage then ensures consistency, coverage, and factual accuracy in generating the final survey. Experiments, including evaluations by human experts and LLM-as-a-judge, demonstrate that SurveyG outperforms state-of-the-art frameworks, producing surveys that are more comprehensive and better structured to the underlying knowledge taxonomy of a field. Our implementation can be found at https://github.com/akBear23/SurveyG.git.
Automated Survey Generation, Literature Synthesis
$c$s: Computing methodologies Information extraction
Figure 1. Standard automated survey pipeline: (1) collect relevant papers, (2) build a structured outline, and (3) write the full survey.
@@ 1. Introduction
The exponential growth of research publications, particularly in rapidly evolving fields such as Artificial Intelligence (Huynh-The et al., 2023), has made it increasingly difficult for researchers to keep pace with new developments (Wang et al., ; Wen et al., 2025). While survey papers serve as invaluable resources by synthesizing existing knowledge and identifying emerging trends, their manual construction is costly, time-consuming, and often unable to keep up with the overwhelming influx of literature (Liang et al., 2025). Although large language models (LLMs) offer promising text generation capabilities, they face critical limitations in handling massive reference sets, maintaining academic rigor, and providing up-to-date knowledge (Wu et al., 2025; Han et al., 2024). These challenges underscore the urgent need for an automated survey generation framework that can efficiently retrieve, organize, and synthesize literature into coherent, high-quality surveys tailored to users’ research interests.
@@
Some recent studies (Su et al., 2025; Wang et al., ; Liang et al., 2025; Yan et al., 2025) have proposed autonomous survey generation frameworks based on user queries, following the basic pipeline illustrated in Figure 1. While these approaches represent promising progress, they exhibit two key limitations. Firstly, they neglect the relationships between papers, such as citation links, methodological connections, or subtopic dependencies, which are essential for understanding how works build upon one another, improve over foundational methods, and collectively shape research trends. Secondly, these frameworks employ a naive strategy for constructing structured outlines or full survey papers, simply concatenating summaries of individual papers. This not only exacerbates the long-context problem in LLMs but also fails to exploit the hierarchical organization of related works within subtopics.
@@
To address these limitations, we propose SurveyG, an autonomous survey generation system that emphasizes knowledge representation and hierarchical summarization. We design an LLM-based multi-agent framework that represents knowledge using a hierarchical citation graph, where nodes correspond to papers and edges capture citation relationships and semantic similarity. The graph is organized into three layers: Foundation, Development, and Frontier, reflecting research progression from seminal contributions to emerging directions. By combining horizontal searches within layers and vertical traversals across layers, our framework generates multi-aspect summaries that are consolidated into structured survey outlines. We evaluate SurveyG on 10 computer science topics from the SurGE benchmark (Su et al., 2025), comparing against state-of-the-art frameworks (Liang et al., 2025; Yan et al., 2025; Wang et al., ) along five dimensions: Coverage, Structure, Relevance, Synthesis, and Critical Analysis. Our key contributions are: (1) a hierarchical citation graph representation modeling both citation and semantic relationships, (2) a graph-based traversal mechanism that produces diverse, multi-aspect summarizations capturing methodological foundations, developmental trends, and frontier directions, and (3) a multi-agent framework combining retrieval-augmented generation with hierarchical summaries as memory to construct coherent, evidence-grounded survey drafts.
@@ 2. Methodology
We introduce SurveyG, an automated survey generation framework with two phases: a preparation phase that retrieves and summarizes papers, builds a hierarchical citation graph, and extracts relationships, and a generation phase that creates a structured outline and composes the full survey through instruction-prompted multi-agent collaboration (Figure 2).
@@ 2.1. Preparation Phase
We represent the relationships among papers using a hierarchical citation graph, where nodes correspond to academic papers and edges capture both citation links and semantic similarity, each weighted by a value $w$. Each node is further assigned to one of three layers: Foundation, Development, or Frontier, which reflect the role of the paper in the progression of research. Formally, the hierarchical citation graph is defined as $G=(V,E,L),$ where $V$ denotes the set of nodes (papers), $E\subseteq V\times V$ is the set of directed or undirected edges encoding citation or semantic relationships, and
$L:V\to\{\text{Foundation},\text{Development},\text{Frontier}\}$
is a layer assignment function. For each node $v_{i}\in V$, we associate a corresponding document $d_{i}\in D$, where $D$ is a database storing the complete content of all papers. In addition, each node $v_{i}\in V$ is equipped with attributes that include a summarization of $d_{i}$ as well as metadata such as the paper’s title and publication year.
@@ 2.1.1. Searching Relevant Paper
where $\text{sim}(\cdot)$ denotes the cosine similarity between the text embedding vectors of the two papers. We consider two embedding strategies: abstract-based embeddings, which are efficient and widely available but may be incomplete or overstated, and LLM-generated summary-based embeddings, which better capture methodological and empirical content but require full text and extra processing (Zhang et al., 2025). We use summary-based embeddings by default because our pipeline already produces content-focused summaries (as detailed in the paragraph below), enabling richer semantic representations of methods, tasks, and domain contributions, and falls back to abstract-based embeddings when full-text access is unavailable.
To better leverage the key content of each paper for survey generation, every node $v_{i}$ is enriched with a summarization derived from its corresponding document $d_{i}$. We design specialized prompt templates tailored to different paper types (Liang et al., 2025) (e.g., surveys, methodological contributions, benchmarks, theoretical works) to extract salient information, including research objectives, methodological approaches, key findings, and limitations. After this phase, we obtain a flat graph $\hat{G}=(V,E)$ that encompasses the papers relevant to the user’s topic along with their relationships. Each node $v_{i}\in V$ is associated with a set of attributes defined as
@@ 2.1.2. Knowledge Representation
To reflect the developmental progression of research within a topic, we assign each node in the flat graph $\hat{G}=(V,E)$ to one of three hierarchical layers via a layer assignment function. (1) Foundation Layer: The foundation layer consists of seminal and high-impact works that form the intellectual backbone of the field. For each paper $p$, we define a trending score as
@@
You are a research analyst synthesizing papers on the topic [QUERY]. <think> Explain your reasoning for clustering papers into 2-3 subgroups based on methodology, contribution, or thematic focus. </think> For each subgroup, summarize the shared methodological approaches, thematic contributions, and provide a concise critique comparing the works. Finally, synthesize an overall perspective highlighting how these subgroups collectively operate in the field.
@@ 2.1.3. Traversal on Graph Strategy
In the vertical stage, we model cross-layer dependencies by tracing the evolution of ideas from each foundation paper. For every seed $s\in V_{\text{foundation}}$, Algorithm 1 performs a weighted breadth-first search (WBFS) over the citation graph, prioritizing semantically relevant nodes via edge weights. The procedure first extracts the foundation paper ($P_{1}$), then collects Development-layer papers reachable from $s$ ($P_{2}$), and finally expands to Frontier-layer papers from $P_{2}$ ($P_{3}$). We summarize these papers hierarchically: an intermediate summary $T_{\text{dev}}$ is generated from $P_{1}\cup P_{2}$, which is then used to incorporate $P_{3}$ and produce the final path summary $T_{\text{path}}$. This staged design leverages temporal progression to reduce long-context issues and improve insight extraction (Zhang et al., 2024). Repeating this process for all $K$ foundation papers yields $K$ vertical summaries, which are combined with $N$ horizontal layer summaries for a total of $K+N$ outputs.
In contrast to earlier frameworks (Wang et al., ; Wen et al., 2025; Liang et al., 2025; Yan et al., 2025) that represent papers as isolated records in a flat database and depend exclusively on RAG-based retrieval, SurveyG organizes the literature within a hierarchical citation graph $G$. This representation integrates both citation and semantic connections among papers, allowing the system to capture the logical progression of research topics over time. By traversing this hierarchy, SurveyG produces a series of summarizations across multiple layers, effectively revealing methodological developments, evolutionary patterns, and current research frontiers. Such a design provides a more coherent and interpretable knowledge foundation for automated survey generation.
Figure 4. Structured outline creation short version prompt.
@@
2.2. Generation Phase
We use a multi-agent conversational framework (Wu et al., 2024) for survey generation, consisting of a Writing Agent with memory initialized from $K+N$ graph-based summaries (Sumers et al., 2023) and an Evaluation Agent that provides diversity-focused feedback using LLM reasoning. Through iterative interaction, the Writing Agent drafts content grounded in summaries, while the Evaluation Agent critiques and refines it for coherence and balance, enabling collaborative survey construction, detailed in Algorithm 2.
@@ 2.2.1. Structured Outline Construction
The Writing Agent constructs an initial structured outline by grounding each section and subsection in the $K+N$ multi-aspect summarizations, ensuring both factual grounding and thematic coherence. The Evaluation Agent then reviews the draft, assessing logical flow and suggesting refinements without altering the overall structure. After the feedback iterations, the outline converges into a coherent and evidence-supported framework. Detailed prompting for both agents is provided in Figure 4 and Appendix D. The key innovation of SurveyG lies in its ability to manage long-context survey synthesis without concatenating all reference texts (Wang et al., ; Liang et al., 2025) or relying on pre-existing human-written surveys (Yan et al., 2025). Instead, it leverages hierarchical summarization from the citation graph $G$ as structured knowledge injected into the Writing Agent.
LLM Model Coverage Structure Relevance Synthesis Critical Analysis
Claude AutoSurvey 73.6(± 3.2)73.6\,($\pm\,$3.2) 64.5(± 4.1)64.5\,($\pm\,$4.1) 80.2(± 2.8)80.2\,($\pm\,$2.8) 51.1(± 5.3)51.1\,($\pm\,$5.3) 45.8(± 4.7)45.8\,($\pm\,$4.7)
SurveyX 74.2(± 3.7)74.2\,($\pm\,$3.7) 71.3(± 4.3)71.3\,($\pm\,$4.3) 82.3(± 3.1)82.3\,($\pm\,$3.1) 62.4(± 4.9)62.4\,($\pm\,$4.9) 52.7(± 5.1)52.7\,($\pm\,$5.1)
SurveyForge 81.8(± 2.9)81.8\,($\pm\,$2.9) 78.4(± 3.5)78.4\,($\pm\,$3.5) 89.1(± 2.1)89.1\,($\pm\,$2.1) 75.4(± 3.8)75.4\,($\pm\,$3.8) 70.3(± 4.2)70.3\,($\pm\,$4.2)
SurveyG ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{88.1}}\,(\pm\,2.4)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{87.9}}\,(\pm\,2.7)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{93.6}}\,(\pm\,1.9)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{80.2}}\,(\pm\,3.2)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{77.3}}\,(\pm\,3.5)$
GPT AutoSurvey 90.7(± 2.1)90.7\,($\pm\,$2.1) 86.2(± 2.8)86.2\,($\pm\,$2.8) 89.3(± 2.5)89.3\,($\pm\,$2.5) 84.6(± 3.1)84.6\,($\pm\,$3.1) 82.3(± 3.4)82.3\,($\pm\,$3.4)
SurveyX 89.3(± 2.6)89.3\,($\pm\,$2.6) 84.4(± 3.2)84.4\,($\pm\,$3.2) 88.5(± 2.8)88.5\,($\pm\,$2.8) 79.2(± 3.9)79.2\,($\pm\,$3.9) 78.4(± 4.1)78.4\,($\pm\,$4.1)
SurveyForge 94.2(± 1.8)94.2\,($\pm\,$1.8) 87.3(± 2.5)87.3\,($\pm\,$2.5) 94.8(± 1.6)94.8\,($\pm\,$1.6) 88.6(± 2.7)88.6\,($\pm\,$2.7) 88.5(± 2.9)88.5\,($\pm\,$2.9)
SurveyG ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{95.7}}\,(\pm\,1.5)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{88.5}}\,(\pm\,2.2)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{95.1}}\,(\pm\,1.4)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{92.2}}\,(\pm\,2.3)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{91.2}}\,(\pm\,2.6)$
Deepseek AutoSurvey 86.5(± 2.7)86.5\,($\pm\,$2.7) 80.4(± 3.6)80.4\,($\pm\,$3.6) 87.8(± 2.9)87.8\,($\pm\,$2.9) 77.3(± 4.2)77.3\,($\pm\,$4.2) 72.4(± 4.5)72.4\,($\pm\,$4.5)
SurveyX 85.4(± 3.1)85.4\,($\pm\,$3.1) 81.2(± 3.8)81.2\,($\pm\,$3.8) 89.4(± 2.7)89.4\,($\pm\,$2.7) 78.2(± 4.3)78.2\,($\pm\,$4.3) 75.6(± 4.6)75.6\,($\pm\,$4.6)
SurveyForge ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{89.4}}\,(\pm\,2.3)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{86.5}}\,(\pm\,2.9)$ 92.5(± 2.0)92.5\,($\pm\,$2.0) 84.1(± 3.4)84.1\,($\pm\,$3.4) 80.7(± 3.7)80.7\,($\pm\,$3.7)
SurveyG 88.7(± 2.5)88.7\,($\pm\,$2.5) 85.7(± 3.0)85.7\,($\pm\,$3.0) ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{94.2}}\,(\pm\,1.8)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{86.3}}\,(\pm\,3.1)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{82.7}}\,(\pm\,3.5)$
Gemini AutoSurvey 94.2(± 1.9)94.2\,($\pm\,$1.9) 70.8(± 5.2)70.8\,($\pm\,$5.2) 96.5(± 1.5)96.5\,($\pm\,$1.5) 83.6(± 3.6)83.6\,($\pm\,$3.6) 84.2(± 3.8)84.2\,($\pm\,$3.8)
SurveyX 90.2(± 2.8)90.2\,($\pm\,$2.8) 71.2(± 5.4)71.2\,($\pm\,$5.4) 95.6(± 1.8)95.6\,($\pm\,$1.8) 82.8(± 3.9)82.8\,($\pm\,$3.9) 84.5(± 4.0)84.5\,($\pm\,$4.0)
SurveyForge 94.9(± 1.7)94.9\,($\pm\,$1.7) 93.1(± 2.1)93.1\,($\pm\,$2.1) 98.7(± 1.1)98.7\,($\pm\,$1.1) 93.5(± 2.4)93.5\,($\pm\,$2.4) 94.2(± 2.5)94.2\,($\pm\,$2.5)
SurveyG ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{96.2}}\,(\pm\,1.4)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{89.4}}\,(\pm\,2.6)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{98.6}}\,(\pm\,1.0)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{95.6}}\,(\pm\,2.1)$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{96.2}}\,(\pm\,2.3)$
Table 1. LLM-as-a-judge evaluation of generated surveys using GPT-4o-mini backbone (mean (± std) over 10 trials per topic). Each LLM evaluates four generation models across content quality dimensions.